# Specialization after Generalization: Towards Understanding Test-Time Training in Foundation Models

### v1

@@ 6 Conclusion
A better understanding of specialization in foundation models opens up several exciting directions for future research. While TTT can be viewed as dynamically specializing a dense model, mixture of experts (MoEs) explicitly increase representational capacity through sparse activation within a larger model. Our findings indicate that investigating the connections of TTT and MoEs could lead to more efficient adaptation strategies and potentially an improved understanding of MoEs. Furthermore, it would be interesting to analyze the compute-efficiency trade-offs of TTT; estimating at which model scale and inference budget TTT becomes beneficial. The importance of specialization is also evident in online learning, where increasingly specialized experience may require ever-larger models to fit globally; TTT offers a potential solution by enabling local adaptation. Particularly interesting would therefore be extending this framework beyond supervised learning to test-time online reinforcement learning, which has recently shown empirical success.
@@ Acknowledgments
We would like to thank Bruce Lee, Celestine Mendler-Dünner, Lars Lorch, and Pierre Bizeul for feedback on early versions of the paper. We also thank Reese Pathak for helpful discussions. JH was supported by the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545. PW was supported by the Max Planck ETH Center for Learning Systems. AS was supported by the Swiss National Science Foundation under grant 204439. GK is partially supported by the Department of Computer Science at ETH Zurich, and the initial part of this work was conducted during his visit to the IDEAL Institute, which was supported by NSF ECCS-2217023.
@@
appendix.Asubsection.A.1subsection.A.2appendix.Bassumption.1assumption.2assumption.3assumption.5appendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4appendix.Dappendix.Esubsection.E.1subsection.E.2subsection.E.2subsection.E.3subsection.E.3equation.E.23subsection.E.4subsection.E.5subsection.E.5subsection.E.5subsection.E.5subsection.E.5subsection.E.5table.caption.19subsection.E.6
@@ A.1 Limitations of global learning under the LRH
We consider the Gaussian measure $\gamma_{d}$ on $\mathbb{R}^{d}$ and define the family of $s$-sparse, $k$-locally linear functions (with $k$ “cells”) denoted by $\mathcal{F}_{d,s,k}$ as the following family:
$\bigg\{f^{\star}:\mathbb{R}^{d}\to\mathbb{R}:$ $x\mapsto\sum_{i=1}^{k}w_{i}^{\top}x\cdot\mathbf{1}_{K_{i}},\forall 1\leq i<j\leq k\;\gamma_{d}(K_{i}\cap K_{j})=0,\|w_{i}\|_{0}\leq s$ (4)
$\|f^{\star}\|_{L_{2}(\gamma_{d})}\asymp 1,\sum_{i=1}^{d}\|\partial_{i}f^{\star}\|_{L_{1}(\gamma_{d})}^{2}\lesssim\frac{s}{d}\bigg\}.$
The final condition defines the sparsity index derived from the $L_{1}$-$L_{2}$ Talagrand’s inequality (cf. the monograph of Chatterjee (2014) and the survey of Ledoux (2019)). This assumption is less restrictive than imposing a bounded Lipschitz constraint. For intuition, consider the function $f^{\star}(x)=\|x\|_{\infty}$; here $k=d$, $s=1$, and the sparsity index condition holds with $\sum_{i=1}^{d}\|\partial_{i}f^{\star}\|_{L_{1}(\gamma_{d})}^{2}\lesssim 1/d$.
We argue that when $s\asymp 1$ and $k$ is polynomial in $d$, the functions in $\mathcal{F}_{d,s,k}$ predominantly lie in the “high” frequencies of the Hermite polynomial basis. Specifically, leveraging results of Paouris et al. (2022), we show that for some constants c,C>0{c,C>0}, it holds that
@@
where $\mathcal{P}_{m}(\cdot)$ denotes the orthogonal projection onto the $m$-degree Hermite polynomial basis.
Geometrically, this spectral concentration implies that most of the gradient’s energy is localized at the decision boundaries between the cells $\{K_{i}\}$. An “underparameterized” global classifier, in this case, the best low-degree approximation, necessarily smooths these boundaries. Therefore, by isoperimetry in high dimensions, this smoothing leads to significant overlap between the decision boundaries of cells. Since most samples fall in the areas of these “distorted/smoothed” decision boundaries, we need many samples to learn such functions (it is well-known that the approximation rate is in $L_{2}$, and the low-degree polynomial is highly non-Lipschitz). The latter aligns with previous observations regarding the required complexity (and the lack of robustness) of foundation models (e.g., Bubeck & Sellke, 2021). We state the main result of this part, whose proof is based on recent works of Paouris et al. (2022) and Damian et al. (2024):
@@
Assume that $\sigma\lesssim 1$, $k\asymp\mathrm{Poly}(d)$, and $s\asymp 1$}. Assume $n\gtrsim\exp(\Omega((\log d)^{2}))$. Then, Equation 5 holds, and there exists a polynomial-time algorithm $\mathcal{A}$ such that
Furthermore, under the low-degree polynomial conjecture, this bound is sharp for several classes of algorithms, and the statistical complexity of any algorithm is at least $\mathrm{Poly}(d)$.
@@ Appendix B Formal assumptions
Next, we formalize the series of hypotheses from Section 5.1.
@@
Even if the global map from concepts $\Phi(x)$ to features $\Psi(x)$ may be non-linear, Assumption 4 posits that a local linear approximation is sufficient within a small neighborhood, where the local behavior can often be approximated by a first-order Taylor expansion. This assumption is further supported by the effectiveness of linear decoders in sparse autoencoders (Elhage et al., 2022).
@@
The feature map $\Psi$ is expressive enough to represent the local function $w_{x^{\star}}$. We consider the set of $s^{\prime}$-sparse weight vectors in $\Phi$-space that are linearly representable by the features $\Psi$. By local linearity (Assumption 4), this means a vector $w$ must lie in the image of $P_{x^{\star}}^{\top}$ (i.e., $w=P_{x^{\star}}^{\top}v$ for some $v\in\mathbb{R}^{d_{2}}$). Let $\tilde{w}_{x^{\star}}$ be the vector in this set that best approximates the predictions of $w_{x^{\star}}$ over the neighborhood. We assume that the resulting representation error is bounded:
@@
If $w_{x^{\star}}$ is not in the image of $P_{x^{\star}}^{\top}$, there is no corresponding vector $v$ in the feature space that can replicate its behavior. This means the compression defined by $P_{x^{\star}}$ has discarded information necessary to represent $w_{x^{\star}}$. Thus, Assumption 5 highlights the importance of pre-training: for the representation error $\eta_{\text{rep}}$ to be small, the feature map needs to learn sufficient structure to represent the ground truth function locally.
@@
Finally, we assume that $\mathbf{\Psi}_{x^{\star}}$ satisfies the generalized restricted eigenvalue (GRE) condition. Combined with local linearity (Assumption 4), the GRE is simply the standard restricted eigenvalue condition on the local concept design matrix $\mathbf{\Phi}_{x^{\star}}$, a condition that is fundamental for guaranteeing stable recovery in sparse regression (Bickel et al., 2009; Van De Geer & Bühlmann, 2009).
@@
Fix any test point $x^{\star}\in\mathcal{X}$ and any $\delta\in(0,1)$. Let Assumptions 1, 2, 3, 4, 5, 6, 7, 8 and 9 hold. Define the inherent misspecification error of any sparse linear approximation as $\smash{\eta_{\text{inherent}}:=\langle\Phi(x^{\star}),w_{\star}-\tilde{w}_{x^{\star}}\rangle^{2}}$ and the misspecification error on the neighborhood as $\smash{\eta_{\text{mis}}:=\eta_{\text{spa}}(r+\eta_{\text{ang}})+\eta_{\text{rep}}}$.
@@
We prove the result in Appendix C and briefly highlight several aspects of the error bound:
@@
    The optimal neighborhood size $k$ depends on the tradeoff between variance (i.e., $\smash{\widetilde{O}(s/k)}$) and bias (from $\eta_{\text{spa}}$). Note that Assumption 3 assumes the neighborhood is described by a $\Theta(s)$-sparse model, which is only possible if $k\ll N$ and the neighborhood is sufficiently “local”.
@@ C.1 Proof of Proposition 2
Fix any test point $x^{\star}\in\mathcal{X}$ and any $\delta\in(0,1)$. Let Assumptions 1, 2, 3, 4, 5, 6, 7, 8 and 9 hold. Let $k$ be the neighborhood size and $s^{\prime}$ the sparsity in concept space. Let concept vectors be bounded in $L_{\infty}$ and $L_{2}$ space with constants $C_{\Phi,\infty},C_{\Phi,2}>0$. Assume the GRE holds with $\kappa>0$. Define the inherent misspecification bias of any sparse linear approximation as $\mathcal{E}_{\text{inherent}}^{2}:=\langle\Phi(x^{\star}),w_{\star}-\tilde{w}_{x^{\star}}\rangle^{2}$.
@@
Step 1: Data decomposition and comparison to oracle. By Assumption 5, there exists some “oracle” $\tilde{v}_{x^{\star}}\in\mathbb{R}^{d_{2}}$ such that $P_{x^{\star}}^{\top}\tilde{v}_{x^{\star}}=\tilde{w}_{x^{\star}}$.
We decompose the observations ${\bm{y}}_{x^{\star}}$ (cf. Assumption 8).
@@
Note that using Assumptions 2, 3 and 5, the average squared magnitude of $\Delta$ is bounded by
@@
Let $h=\tilde{v}_{x^{\star}}-\hat{v}_{x^{\star}}^{\text{TTT}}$. Since by Assumption 5, both $\tilde{v}_{x^{\star}}$ and $\hat{v}_{x^{\star}}^{\text{TTT}}$ satisfy the constraint of Equation 6, the error vector $h$ is sparse in the concept space: $\|P_{x^{\star}}^{\top}h\|_{0}\leq 2s^{\prime}$.
Step 2: By the optimality of $\hat{v}_{x^{\star}}^{\text{TTT}}$ (cf. Assumption 6):
@@
Substituting ${\bm{y}}_{x^{\star}}=\mathbf{\Psi}_{x^{\star}}\tilde{v}_{x^{\star}}+\Delta+\varepsilon$ (using local linearity, Assumption 4, $\mathbf{\Psi}_{x^{\star}}\tilde{v}_{x^{\star}}=\mathbf{\Phi}_{x^{\star}}\tilde{w}_{x^{\star}}$):
@@
Step 3: We analyze the right hand side of Equation 9. Using local linearity (cf. Assumption 4), $h^{\top}\mathbf{\Psi}_{x^{\star}}^{\top}=(P_{x^{\star}}^{\top}h)^{\top}\mathbf{\Phi}_{x^{\star}}^{\top}$. Let $Z=\frac{1}{k}\mathbf{\Phi}_{x^{\star}}^{\top}(\Delta+\varepsilon)$ be the total concept score. Then,
@@
Substituting this back into Equation 9 gives:
@@
Step 4: Applying GRE. Since $\|P_{x^{\star}}^{\top}h\|_{0}\leq 2s^{\prime}$, we can apply the GRE (cf. Assumption 9) to Equation 10:
@@
We next bound the estimation error $\mathcal{E}_{\text{estimation}}$ using Assumption 7 ($L_{2}$ bound):
@@
Combining this with Equation 11 gives:
@@ C.2 Concentration bounds
Under Assumption 7 ($L_{\infty}$ bound) and Assumption 8 (subgaussian noise), there exists a universal constant $C_{H}>0$ such that for any $\delta\in(0,1)$, with probability at least $1-\delta$:
@@
Let $\Delta$ be a misspecification vector such that $\frac{1}{k}\|\Delta\|_{2}^{2}\leq\eta$. Under Assumption 7 ($L_{\infty}$ bound), the correlation of $\Delta$ with sparse concept vectors is bounded by:

### v2

@@ 6 Conclusion
A better understanding of specialization in foundation models opens up several exciting directions for future research. An interesting question is understanding what determines the optimal neighborhood size and whether it depends on the test point. Furthermore, it would be interesting to analyze the compute-efficiency trade-offs of TTT; estimating at which model scale and inference budget TTT becomes beneficial. The importance of specialization is also evident in online learning, where increasingly specialized experience may require ever-larger models to fit globally; TTT offers a potential solution by enabling local adaptation. Particularly interesting would therefore be extending this framework beyond supervised learning to test-time online reinforcement learning, which has recently shown empirical success.
@@ Acknowledgments
We would like to thank Bruce Lee, Celestine Mendler-Dünner, and Lars Lorch for feedback on early versions of the paper. We also thank Reese Pathak and Pierre Bizeul for helpful discussions. JH was supported by the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545. PW was supported by the Max Planck ETH Center for Learning Systems. AS was supported by the Swiss National Science Foundation under grant 204439. GK conducted the initial part of this work during his visit to the IDEAL Institute, hosted by Lev Reyzin, which was supported by NSF ECCS-2217023.
@@
appendix.Asubsection.A.1subsection.A.2appendix.Bassumption.1assumption.2assumption.3assumption.5appendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4appendix.Dappendix.Esubsection.E.1subsection.E.2subsection.E.2subsection.E.3subsection.E.3equation.E.23subsection.E.4subsection.E.5subsection.E.6subsection.E.6subsection.E.6subsection.E.6subsection.E.6subsection.E.6table.caption.19subsection.E.7
@@ A.1 Limitations of global learning under the LRH
We consider the Gaussian measure $\gamma_{d}$ on $\mathbb{R}^{d}$ and define the family of $s$-sparse, $k$-locally linear functions (with $k$ “cells”), denoted by $\mathcal{F}_{d,s,k}\subset\{\mathbb{R}^{d}\to\mathbb{R}\}$, as defined as
$\mathcal{F}_{d,s,k}=\bigg\{f^{\star}(x)=\sum_{i=1}^{k}w_{i}^{\top}x\cdot\mathbf{1}_{K_{i}}:\ \forall$ $1\leq i<j\leq k\;\gamma_{d}(K_{i}\cap K_{j})=0,\|w_{i}\|_{0}\leq s,\|w_{i}\|_{2}\lesssim 1$ (4)
$\|f^{\star}\|_{L_{2}(\gamma_{d})}\asymp 1,\sum_{i=1}^{d}\|\partial_{i}f^{\star}\|_{L_{1}(\gamma_{d})}^{2}\lesssim\frac{s}{d}\bigg\},.$
The final condition defines the sparsity index derived from the $L_{1}$-$L_{2}$ Talagrand’s inequality (cf. the monograph of Chatterjee (2014) and the survey of Ledoux (2019)), and note that the cells $K_{1},\ldots,K_{k}$ may depend on the function $f^{\star}$. For intuition, consider the function $f^{\star}(x)=\|x\|_{\infty}$; here $k=d$, $s=1$, and the sparsity index condition holds with $\sum_{i=1}^{d}\|\partial_{i}f^{\star}\|_{L_{1}(\gamma_{d})}^{2}\lesssim 1/d$.
We argue that when $s\asymp 1$ and $k$ is polynomial in $d$, the functions in $\mathcal{F}_{d,s,k}$ predominantly lie in the “high” frequencies of the Hermite polynomial basis. Specifically, by leveraging results of Paouris et al. (2022), we show that for some constants c,C>0{c,C>0}, it holds that
@@
where $\mathcal{P}_{m}(\cdot)$ denotes the orthogonal projection onto the $m$-degree Hermite polynomial basis. The main result of this part is the following:
@@
Assume that $\sigma\lesssim 1$, $k\asymp\mathrm{Poly}(d)$, and $s\asymp 1$}. Then Equation 5 holds. Furthermore, for $n\gtrsim\exp(\Omega((\log d)^{2}))$, there exists a polynomial-time algorithm $\mathcal{A}$ such that
$\sup_{f^{\star}\in\mathcal{F}_{d,s,k}}\mathbb{E}_{\mathcal{D}}\|\mathcal{A}(\mathcal{D})-f^{\star}\|_{L_{2}(\gamma_{d})}\leq 0.1,$
and under the low-degree polynomial conjecture, this bound is sharp for few classes of algorithms.
We refer to the survey of Reyzin (2020) for more details on the low-degree polynomial conjecture. We emphasize that our assumptions on $\mathcal{F}_{d,s,k}$ are much less restrictive than imposing a uniform bounded Lipschitz (or an average $L_{2}$ Lipschitz constraint), as the Lipschitz constant may be high on the boundary between two cells, and the results of Bizeul & Klartag (2025) implying that Lipschitz functions can be learned with polynomial number of samples.
Roughly speaking, from a geometric view, this spectral concentration implies that most of the energy of the coefficients is localized at the decision boundaries between the cells $\{K_{i}\}$. An “underparameterized” global classifier, in this case, the best low-degree approximation, necessarily smooths these boundaries. Therefore, by isoperimetry in high dimensions, this smoothing leads to significant overlap between the decision boundaries of cells. Since most samples fall in the areas of these “distorted/smoothed” decision boundaries, we need many samples to learn such functions. [^8] It is well-known that even for $1$-Lipschitz function, its best low-degree polynomial (it terms of $L_{2}$), is highly non-Lipschitz cf. Bizeul & Klartag (2025). The latter aligns with previous observations regarding the required complexity (and the lack of robustness) of foundation models (e.g., Bubeck & Sellke, 2021).
@@ Appendix B Formal assumptions
Next, we formalize the series of hypotheses from Section˜5.1.
@@
Even if the global map from concepts $\Phi(x)$ to features $\Psi(x)$ may be non-linear, Assumption˜4 posits that a local linear approximation is sufficient within a small neighborhood, where the local behavior can often be approximated by a first-order Taylor expansion. This assumption is further supported by the effectiveness of linear decoders in sparse autoencoders (Elhage et al., 2022).
@@
The feature map $\Psi$ is expressive enough to represent the local function $w_{x^{\star}}$. We consider the set of $s^{\prime}$-sparse weight vectors in $\Phi$-space that are linearly representable by the features $\Psi$. By local linearity (Assumption˜4), this means a vector $w$ must lie in the image of $P_{x^{\star}}^{\top}$ (i.e., $w=P_{x^{\star}}^{\top}v$ for some $v\in\mathbb{R}^{d_{2}}$). Let $\tilde{w}_{x^{\star}}$ be the vector in this set that best approximates the predictions of $w_{x^{\star}}$ over the neighborhood. We assume that the resulting representation error is bounded:
@@
If $w_{x^{\star}}$ is not in the image of $P_{x^{\star}}^{\top}$, there is no corresponding vector $v$ in the feature space that can replicate its behavior. This means the compression defined by $P_{x^{\star}}$ has discarded information necessary to represent $w_{x^{\star}}$. Thus, Assumption˜5 highlights the importance of pre-training: for the representation error $\eta_{\text{rep}}$ to be small, the feature map needs to learn sufficient structure to represent the ground truth function locally.
@@
Finally, we assume that $\mathbf{\Psi}_{x^{\star}}$ satisfies the generalized restricted eigenvalue (GRE) condition. Combined with local linearity (Assumption˜4), the GRE is simply the standard restricted eigenvalue condition on the local concept design matrix $\mathbf{\Phi}_{x^{\star}}$, a condition that is fundamental for guaranteeing stable recovery in sparse regression (Bickel et al., 2009; Van De Geer & Bühlmann, 2009).
@@
Fix any test point $x^{\star}\in\mathcal{X}$ and any $\delta\in(0,1)$. Let Assumptions˜1, 2, 3, 4, 5, 6, 7, 8 and 9 hold. Define the inherent misspecification error of any sparse linear approximation as $\smash{\eta_{\text{inherent}}:=\langle\Phi(x^{\star}),w_{\star}-\tilde{w}_{x^{\star}}\rangle^{2}}$ and the misspecification error on the neighborhood as $\smash{\eta_{\text{mis}}:=\eta_{\text{spa}}(r+\eta_{\text{ang}})+\eta_{\text{rep}}}$.
@@
We prove the result in Appendix˜C and briefly highlight several aspects of the error bound:
@@
    The optimal neighborhood size $k$ depends on the tradeoff between variance (i.e., $\smash{\widetilde{O}(s/k)}$) and bias (from $\eta_{\text{spa}}$). Note that Assumption˜3 assumes the neighborhood is described by a $\Theta(s)$-sparse model, which is only possible if $k\ll N$ and the neighborhood is sufficiently “local”.
@@ C.1 Proof of Proposition 2
Fix any test point $x^{\star}\in\mathcal{X}$ and any $\delta\in(0,1)$. Let Assumptions˜1, 2, 3, 4, 5, 6, 7, 8 and 9 hold. Let $k$ be the neighborhood size and $s^{\prime}$ the sparsity in concept space. Let concept vectors be bounded in $L_{\infty}$ and $L_{2}$ space with constants $C_{\Phi,\infty},C_{\Phi,2}>0$. Assume the GRE holds with $\kappa>0$. Define the inherent misspecification bias of any sparse linear approximation as $\mathcal{E}_{\text{inherent}}^{2}:=\langle\Phi(x^{\star}),w_{\star}-\tilde{w}_{x^{\star}}\rangle^{2}$.
@@
Step 1: Data decomposition and comparison to oracle. By Assumption˜5, there exists some “oracle” $\tilde{v}_{x^{\star}}\in\mathbb{R}^{d_{2}}$ such that $P_{x^{\star}}^{\top}\tilde{v}_{x^{\star}}=\tilde{w}_{x^{\star}}$.
We decompose the observations ${\bm{y}}_{x^{\star}}$ (cf. Assumption˜8).
@@
Note that using Assumptions˜2, 3 and 5, the average squared magnitude of $\Delta$ is bounded by
@@
Let $h=\tilde{v}_{x^{\star}}-\hat{v}_{x^{\star}}^{\text{TTT}}$. Since by Assumption˜5, both $\tilde{v}_{x^{\star}}$ and $\hat{v}_{x^{\star}}^{\text{TTT}}$ satisfy the constraint of Equation˜6, the error vector $h$ is sparse in the concept space: $\|P_{x^{\star}}^{\top}h\|_{0}\leq 2s^{\prime}$.
Step 2: By the optimality of $\hat{v}_{x^{\star}}^{\text{TTT}}$ (cf. Assumption˜6):
@@
Substituting ${\bm{y}}_{x^{\star}}=\mathbf{\Psi}_{x^{\star}}\tilde{v}_{x^{\star}}+\Delta+\varepsilon$ (using local linearity, Assumption˜4, $\mathbf{\Psi}_{x^{\star}}\tilde{v}_{x^{\star}}=\mathbf{\Phi}_{x^{\star}}\tilde{w}_{x^{\star}}$):
@@
Step 3: We analyze the right hand side of Equation˜9. Using local linearity (cf. Assumption˜4), $h^{\top}\mathbf{\Psi}_{x^{\star}}^{\top}=(P_{x^{\star}}^{\top}h)^{\top}\mathbf{\Phi}_{x^{\star}}^{\top}$. Let $Z=\frac{1}{k}\mathbf{\Phi}_{x^{\star}}^{\top}(\Delta+\varepsilon)$ be the total concept score. Then,
@@
Substituting this back into Equation˜9 gives:
@@
Step 4: Applying GRE. Since $\|P_{x^{\star}}^{\top}h\|_{0}\leq 2s^{\prime}$, we can apply the GRE (cf. Assumption˜9) to Equation˜10:
@@
We next bound the estimation error $\mathcal{E}_{\text{estimation}}$ using Assumption˜7 ($L_{2}$ bound):
@@
Combining this with Equation˜11 gives:
@@ C.2 Concentration bounds
Under Assumption˜7 ($L_{\infty}$ bound) and Assumption˜8 (subgaussian noise), there exists a universal constant $C_{H}>0$ such that for any $\delta\in(0,1)$, with probability at least $1-\delta$:
@@
Let $\Delta$ be a misspecification vector such that $\frac{1}{k}\|\Delta\|_{2}^{2}\leq\eta$. Under Assumption˜7 ($L_{\infty}$ bound), the correlation of $\Delta$ with sparse concept vectors is bounded by: